OpenStudy (anonymous):

8x - 12= 4x please help...

OpenStudy (atlas):

8x - 12 =4x Subtracting 4x from both sides: (because subtracting equal numbers from both sides does not change the equality)

OpenStudy (anonymous):

8x -12 = 4x -4x -4x 4x -12 = 0 +12 +12 4x / 4 = 12 /4 x = 3

OpenStudy (atlas):

8x-12-4x = 4x-4x; or, 4x-12 =0

OpenStudy (texaschic101):

we want to get all the x terms on one side and everything else on the other side. 8x - 12 = 4x --- subtract 8x from both sides 8x - 8x - 12 = 4x - 8x -- combine like terms -12 = - 4x -- now divide both sides by -4 -12/-4 = x 3 = x check.. 8x - 12 = 4x 8(3) - 12 = 4(3) 24 - 12 = 12 12 = 12 (correct) x = 3

OpenStudy (atlas):

or, 4x = 12; Dividing 4 from both sides: (dividing equal number from both sides does not change equality) x =12/4 =3

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

yup